OpenStudy (anonymous):

whats the deritives of 1/x^5?

OpenStudy (anonymous):

\[1/x^5\]

OpenStudy (anonymous):

-5(1/x^6)

OpenStudy (anonymous):

is it 5/x^6?

OpenStudy (anonymous):

wait why negative 5?

OpenStudy (anonymous):

1/x^5=x^-5 right do we agree up to here, now apply the x^k rule=-5x^-6=-5/x^6

OpenStudy (anonymous):

o ok thanks

OpenStudy (anonymous):

remeber 1/x^k will always be -kx^(-k-1)